Ammonia can be synthesized according to the equation:

N2(g)+2H2(g)<--->2NH3(g)

The Kp for this reaction at 725 K is 7.25x10-5; Calculate Kc at this temperature.
Answer #1

ANSWER:

N2(g)+2H2(g)<--->2NH3(g)

The problem can be solved by the following equation:

Kp = Kc (RT)Δn

Δn = (total gaseous products) - (total gaseous reactants) = (2 - 3) = -1

R = gas constant = 8.314 J/Mol K, ;;;; T = Absolute temperature

Kc = Kp / (RT)Δn = 7.25x10-5 / (8.314 X 725) -1 = 7.25x10-5 X (8.314 X 725) +1= 0.44